# Unveiling the electronic structure of the charge density wave and topological semimetal TaTe4 through high-field magnetotransport measurements

### v1

@@ Abstract
Understanding the interplay between topology and correlated electron states is central to the study of quantum materials. TaTe4, a charge density wave (CDW) compound predicted to host topological phases, offers a platform to explore this interplay. Here, we report a comprehensive study of the Fermi surface (FS) of TaTe4 via high-field magnetotransport measurements and density functional theory (DFT) calculations. Using three distinct field-current configurations, we resolve four out of five FS pockets predicted in the CDW phase, with no evidence of residual non-CDW features – in contrast to recent ARPES studies. Remarkably, we identify a previously unobserved quasi-cylindrical hole pocket and a high-frequency quantum oscillation contribution to the magnetoresistance attributable to magnetic breakdown between reconstructed FS sheets, from which we estimate a CDW gap of $\sim$0.29 eV. Additionally, we observe a robust linear magnetoresistance (MR) for all field orientations when current flows along the $a-$axis, with a distinct high-field linear MR regime emerging near the $c-$axis, consistent with FS hot spots and magnetic breakdown. Our results establish TaTe4 as a paradigmatic system where CDW-driven FS reconstruction governs the bulk electronic structure, enabling new routes to probe the novel physics that can arise in the intersection between correlated electronic states and topological band features.
@@ I Introduction
Transition metal tetrachalcogenides (TMTs), i.e. TaTe4 and NbTe4, are promising in this context due to their complex band structures which include Dirac points, and a transition to a charge density wave (CDW) phase that is expected to deeply reconstruct its band structure and leads to the emergence of new high-degeneracy Dirac points [13, 14, 15, 16]. Therefore, these materials stand out as ideal systems for probing the interplay between topological features and CDW order.
@@
In this paper, we present a detailed band structure analysis of TaTe4 using a combination of DFT calculations and high-field magnetoresistance measurements, providing detailed insights into the FS, not addressed in previous works. We report magnetoresistance data for TaTe4 under magnetic fields up to 35 T oriented and rotated along two perpendicular planes, offering a clear image of the Fermi surface of the material. We found that the bulk Fermi surface of this material fully matches the DFT predictions for the CDW-phase with four of five predicted bands identified. No evidence of non-CDW FS was found. In addition, we show that a linear magnetoresistance appears at all angles at low magnetic fields when imposing a current perpendicular to the 1D chains (this is, along the $a$-axis). Furthermore, we report the presence of a high Shubnikov-de Haas (SdH) frequency in the magnetoresistance data at certain angles in this configuration, which is consistent with magnetic breakdown effect,and that a secondary linear magnetoresistance regime appears at high magnetic fields at angles for which magnetic breakdown occurs. Our work provides complete and detailed information of the band structure of this material at the Fermi level, which is paramount in the understanding of the possible connection between topological and correlated phases of matter.
II Materials and methods
Figure 1: (Color online) (a) Unit cell of TaTe4 in both the CDW (large rectangular prism) and non-CDW (small rectangular prism) phases. (b) High symmetry points of the first brillouin zone of TaTe4, in both its high symmetry non-CDW phase (corners of red polygon with capital letters) and low symmetry CDW phase (corners of red polygon with lowercase letters). (c-d) and (e-f) are the band structure and Fermi surface of TaTe4 obtained through DFT calculations for the non-CDW phase and the CDW phase, respectively. In (c,e) bands crossing the Fermi level are highlighted in colors other than black. (g) Four of the five CDW FS pockets shown individually to better illustrate the possible closed orbits for each. The fifth FS pocket, centered around $a$ and $m$ points, is too small to be visualized.
II.1 Computational details of the electronic structure calculations
[…]
II.2 Sample growth
II.3 Sample characterization and transport measurements
III Results and discussion
III.1 Band structure calculations
@@
Once the temperature is lowered below about 400 K, the $P4/mcc$ to the $P4/ncc$ displacive phase transition takes place and the CDW phase is obtained [31], with a new unit cell $2a\times 2a\times 3c$ (large rectangular prism in Fig. 1(a)) due to a A-B-A stacking between adjacent chains. Its new Brillouin zone high symmetry points are shown in the corners of the blue polyhedron of Fig. 1(b), and labeled in lowercase letters. With this, the band structure is drastically modified, opening gaps at the points where the non-CDW Dirac points appeared and leading to the emergence of new high-degeneracy points, including an eightfold-degenerate crossing at point $a$ which has been indexed as a Double Dirac point in previous works [15]. Five bands cross the fermi level (Fig. 1(e)), leading to the FS portrayed at Fig. 1(f). It consists of two electron-like semi-ellipsoidal pockets centered at the x-point of the new Brillouin zone, labeled as $\beta$ and $\epsilon$ in Fig. 1(g), in addition to a couple of hole-like pockets with cylindrical sections, $\alpha$ and $\zeta$. While the $\alpha$ hole-like pocket sections extends along the $z$-$r$-$a$ plane and contains a combination of closed and open orbits, the $\zeta$ hole-like pocket is completely cylindrical, centered at $x$ and extending along the $x$-$r$ direction. A fifth band crossing the Fermi level can be observed, leading to very small electron orbits centered at $m$ and $a$ high symmetry points. However, due to its small size and the limited $k$-space grid used for the interpolation, it is difficult to visualize it. Interestingly, this pocket is also not observed in our magnetoresistance measurements, suggesting that its small size is indeed real.
III.2 Magnetoresistance measurements
III.2.1 Measurement configurations
@@
III.2.2 Oscillatory component of the magnetoresistance
In order to retrieve information from the FS of TaTe4, magnetoresistance measurements were processed to isolate their oscillatory component, which arises from the SdH effect. This was achieved by substracting a smooth background through a spline fit, yielding the oscillatory traces in the plots shown in Figs. 2(a,b,c)(iii). After that, a Fast Fourier transform (FFT) was applied to the oscillatory signal vs inverse field data, and the dominant frequencies were retrieved for all magnetic field orientations in each configuration, as presented in Figs. 2(a,b,c)(iv)). From these frequencies the extremal cross-sectional areas of the Fermi surface ($A_{FS}$) of TaTe4 were determined following the Onsager relation $F_{1/B}=(\phi_{0}/2\pi^{2})A_{FS}$, where ϕ0=subscriptitalic-$\phi_{0}=$2.07$\times$10-15 T m2 is the magnetic flux quantum. Due to the presence of multiple oscillation frequencies, each corresponding to different sections or pockets of the FS of the material, the analysis of the evolution of such frequencies with the magnetic field orientation was conducted using the color map plots of Figs. 3(a,b,c)(i) and (ii). In those plots, the colors represent how dominant each frequency is. Intense yellow indicates dominant frequencies, while dark blue represent frequencies that do not contribute to the oscillatory signal. In order to facilitate the analysis, the frequency space was divided in three different regimes: a high frequency regime, from 1000 T to 5000 T, a mid-frequency regime, from 300 T to 1000 T, and a low frequency regime, below 300 T.
The analysis for different frequency regimes was motivated by recent ARPES measurements [19, 16, 17]. These studies revealed that the most intense spectral weight at the Fermi level are ellipsoids centered at the X/M-point in reciprocal space which follow the periodicity of the high-symmetry non-CDW phase, with a cross sectional area about  0.4 Å$\AA^{-2}$. According to the Onsager relation, such pocket would correspond to a frequency close to 4000 T when field is perpendicular to the a-c plane, which corresponds to 0° and 90° in configuration A (Figs. 3(a)(i,ii)) and 0° for configurations B and C (Figs. 4(a)(i,ii)). However, no frequency component is observed around the a-direction in the high frequency regime in any configuration. Therefore, there is no evidence of such pocket from the SdH measurements. In configuration A and for this high frequency range, the only contribution occurs in a limited range of field orientations close to 45°, where frequencies of 1387 T $\pm$ 8 T, 2774 T $\pm$ 11 T and 4163 T $\pm$ 15 T are detected (see Fig. 3 (a)(i)). These contributions correspond to the fundamental, first, and second harmonics of the 1387 T signal. Although this frequency value lies in the range of closed orbits that could originate in the FS of the non-CDW phase, the angle dependence of this signal does not agree with that of the predicted non-CDW pockets, which are displayed in Fig. 3(a)(iii).
@@ III.2.2 Oscillatory component of the magnetoresistance
Since $\beta$ and $\epsilon$ frequencies are the most dominant in the FFT spectrum, it was possible to track their evolution with temperature and calculate their effective masses. To estimate the contribution of charge carriers to the electronic properties of TaTe4, the Liftshitz-Kosevich (LK) formula for the oscillatory magnetoresistance $\tilde{\rho}(B)$, as presented in equation A1 of the appendix A section, is used [32]. By fitting the evolution of the oscillatory magnetoresistance with temperature as predicted by this equation, the effective cyclotron masses of each pocket for a fixed angle, $m^{cyc}_{\theta}$, can be obtained. In our case, the effective masses were estimated by filtering the oscillatory signal of each pocket with a band-pass filter to select a single frequency, resulting in a better fit to our data than by performing the fit on the sum of all frequencies with the whole LK formula.
Using this approach and following the magnetic field orientations defined for configuration A in previous sections, we obtained $m^{cyc}_{\beta-0^{\circ}}=(0.30\pm 0.03)m_{e}$ for the lower frequency branch of pocket $\beta$ at 0 and $m^{cyc}_{\beta-45^{\circ}}=(0.32\pm 0.03)m_{e}$ at 45 in configuration A (Figs. 6(a,b), and Appendix A for more details), which suggests that the mass of charge carriers in these pockets is highly isotropic along this rotation plane.
The angular evolution of low-frequency oscillations in configuration A is shown in Figs. 3(c)(i,ii). A clear dominant frequency branch is observed, evolving from 47 T $\pm$2 T at 0° and 90° to 30 T±$\pm$3 T around 45°, as highlighted by the red dots in Fig.3(c)(ii). This branch remains visible up to the fourth harmonic. Its angular dependence and magnitude are consistent with the cylindrical sections of the $\alpha$ pocket predicted for the CDW-phase Fermi surface (see Figs. 1(g) and 3(c)(iii)), as emphasized by its comparison to the predicted evolution of this orbit (lower red solid line in Fig. 3(c)(ii)). The minimum frequency occurring around 45° supports the prediction that these sections are elongated cylinders aligned along the z-a direction. A second contribution, arising from the wider sections of the same FS pocket centered at the a-point of the CDW Brillouin zone appears near 150 T at 0° and increases rapidly as the magnetic field is rotated toward 45°. This behavior suggests that both the elongated and wider sections are connected, forming a single, merged FS pocket, as expected from our DFT calculations.
@@
The $\alpha$ pocket is anisotropic in terms of effective mass and shape. Its charge carriers effective mass smoothly changes with angle, with $m^{cyc}_{\alpha-0^{\circ}}=(0.21\pm 0.03)m_{e}$ at 0 and $m^{cycl}_{\alpha-45^{\circ}}=(0.10\pm 0.02)m_{e}$ at 45 in configuration A. Further details of the effective mass analysis can be found in Appendix A.
@@
In contrast to configuration B, configuration C, for which the magnetic field is always perpendicular to the current, exhibited a clear evolution of different dominant oscillation frequencies at all angles in the $a-c$ plane, as shown in Figs. 4(a)(i) and (b)(i). In this case, we defined two different frequency regimes for the analysis: a low/mid frequency regime below 2000 T and a high frequency regime, from 2000 T to 5000 T. As in configuration A, the low/mid-frequency regimes dominated the oscillations in configuration C (Fig. 4(a)(i)). The evolution of the $\beta$ pocket, already clearly identified in configuration A, could be tracked at angles near the a-axis; however, its frequency gradually increases and eventually disappears as the magnetic field orientation approaches the c-axis — an effect also observed in configuration B. In contrast, as the $\beta$ pocket oscillation faded, another frequency of oscillation appeared in the range of 1040 T $\pm$ 8 T at 35° and 829 T $\pm$ 7 T at 90°. This frequency follows a clear cylinder-like angular dependence following $1/\cos(\pi/4-\theta)$, in agreement with the DFT predicted $\zeta$ pocket (Figs. 4(a)(ii,iii)). This feature had not been reported in previous magnetotransport studies, likely due to the experimental challenges of imposing current along the a-axis in TaTe4 derived from the quasi-1D morphology of the crystals, as discussed earlier.
III.2.3 High frequency regime and magnetic breakdown effects
@@
The results discussed in the previous two subsections have provided evidence of four out of the five predicted pockets for the CDW-phase FS, with no evidence of contributions coming from bands remanent from the non-CDW high symmetry phase, in contrast to results by recent ARPES works [17, 16, 18]. We can therefore say that the bulk transport band structure of this material is dominated by the CDW-reconstructed bands, in close proximity of size and shape to the DFT predicted FS.
III.2.4 Linear magnetoresistance
@@
III.3 Conclusions and general remarks
In summary, we reported detailed evidence of the presence of four of five FS pockets predicted by DFT calculations of TaTe4 in the CDW-phase using high-field magnetotransport measurements. The dominant features identified consisted of two ellipsoidal electron pockets ($\beta$ and $\varepsilon$) centered at the $x$ point; a quasi-cylindrical hole-like pocket ($\zeta$) centered at $x$ and with its main axis elongated along the $x-r$ direction, not observed before by other studies; and a hole pocket ($\alpha$) with open and closed orbits, including cylindrical sections with the main axis elongated along the $z-a$ direction. These results indicate that the bulk FS of this compound is fully dominated by the CDW reconstructed FS, with no vestiges of the non-CDW bands. Our results are significant as they provide a firm standing point for further studies on both, the non-trivial topological character of this compound, and the presence of non-trivial topological quasiparticles connected to the CDW state.
In addition, we observed a frequency of oscillation in the SdH signal that is consistent with magnetic breakdown effect between $\alpha-$ and $\zeta$-pockets, from which we could estimate a CDW gap value of 0.29 eV, in accordance to ARPES measurements and band-structure calculations results. Such magnetic breakdown effect could be the origin of the linear magnetoresistance observed above 8 T and at high fields, for magnetic fields oriented closed to the $c-$direction. We also observed a linear magnetoresistance dependece at low fields, with a different slope than in the high-field regime, which is present for all field orientations when current is along the $a-$direction. Although this effect could be related to the topological character of the band structure of TaTe4, further measurements and studies are required.

### v2

# Fermi Surface Reconstruction and Anisotropic Linear Magnetoresistance in the Charge Density Wave Topological Semimetal TaTe4

@@ Abstract
Understanding the interplay between topology and correlated electron states is central to the study of quantum materials. TaTe4 is a quasi-one-dimensional charge density wave (CDW) compound predicted to host topological phases, which makes it a model platform to explore this interplay. Here, we combine high-field magnetotransport measurements with density functional theory calculations to provide a comprehensive mapping of the Fermi surface (FS) of TaTe4 in its CDW phase. Using multiple current-field geometries, we resolve the four largest of six pockets of the FS predicted by theory and find no evidence of non-CDW bands, highlighting the full reconstruction of the FS in the bulk. We identify a previously unobserved quasi-cylindrical pocket and uncover a large size orbit consistent with magnetic breakdown between reconstructed FS sheets, from which we estimate a CDW gap of $\sim$0.29 eV. Moreover, we observe a robust linear magnetoresistance that persists across all field directions when current flows perpendicular to the 1D chains along which the CDW is formed, with a distinct high-field linear regime emerging when field is along the chains. These findings establish TaTe4 as a prototypical material to study the coexistence of correlation-driven reconstruction and topological electronic states.
@@ I Introduction
Transition metal tetrachalcogenides (TMTs), i.e. TaTe4 and NbTe4, are promising in this context due to their complex band structures which include Dirac points, and a transition to a charge density wave (CDW) phase that is expected to deeply reconstruct its band structure and leads to the emergence of new high-degeneracy Dirac points [13, 14, 15, 16, 17]. Therefore, these materials stand out as ideal systems for probing the interplay between topological features and CDW order.
@@
In this paper, we present a detailed band structure analysis of TaTe4 using a combination of DFT calculations and high-field magnetoresistance measurements, providing detailed insights into the FS, not addressed in previous works. We report magnetoresistance data for TaTe4 under magnetic fields up to 35 T oriented and rotated along two perpendicular planes, offering a clear image of the Fermi surface of the material. We found that the bulk Fermi surface of this material fully matches the DFT predictions for the CDW-phase with four of six predicted bands identified. No evidence of non-CDW FS was found. In addition, we show that a linear magnetoresistance appears at all angles at low magnetic fields when imposing a current perpendicular to the 1D chains (this is, along the $a$-axis). Furthermore, we report the presence of a high Shubnikov-de Haas (SdH) frequency in the magnetoresistance data at certain angles in this configuration, which is consistent with magnetic breakdown effect,and that a secondary linear magnetoresistance regime appears at high magnetic fields at angles for which magnetic breakdown occurs. Our work provides complete and detailed information of the band structure of this material at the Fermi level, which is paramount in the understanding of the possible connection between topological and correlated phases of matter.
Figure 1: (Color online) (a) Unit cell of TaTe4 in both the CDW (large rectangular prism) and non-CDW (small rectangular prism) phases. (b) High symmetry points of the first brillouin zone of TaTe4, in both its high symmetry non-CDW phase (corners of red polygon with capital letters) and low symmetry CDW phase (corners of red polygon with lowercase letters). (c-d) and (e-f) are the band structure and Fermi surface of TaTe4 obtained through DFT calculations for the non-CDW phase and the CDW phase, respectively. In (c,e) bands crossing the Fermi level are highlighted in colors other than black. (g) Four of the CDW FS pockets shown individually and in the corresponding band color of (e) to better illustrate the possible closed orbits for each. The other two FS pockets coming from bands in yellow and orange in (e), and centered around $a$ and $m$ points, are too small to be visualized.
II Results and discussion
II.1 Band structure calculations
@@
Once the temperature is lowered below about 400 K, the $P4/mcc$ to the $P4/ncc$ displacive phase transition takes place and the CDW phase is obtained [24], with a new unit cell $2a\times 2a\times 3c$ (large rectangular prism in Fig. 1(a)) due to a A-B-A stacking between adjacent chains. Its new Brillouin zone high symmetry points are shown in the corners of the blue polyhedron of Fig. 1(b), and labeled in lowercase letters. With this, the band structure is drastically modified, opening gaps at the points where the non-CDW Dirac points appeared and leading to the emergence of new high-degeneracy points, including an eightfold-degenerate crossing at point $a$ which has been indexed as a Double Dirac point in previous works [15]. Six bands cross the fermi level (Fig. 1(e)), leading to the FS portrayed at Fig. 1(f). It consists of one hole-like band labeled as $\alpha$ in Fig. 1(g), which extends along the $z$-$r$-$a$ plane of the CDW Brilloiun zone and contains a combination of closed and open orbits; two electron-like semi-ellipsoidal pockets centered at the x-point, labeled as $\beta$ and $\epsilon$; and one electron-like cylindrical pocket dubbed as $\zeta$, with its main axis centered at $x$ and extending along the $x$-$r$ direction. Two additional bands cross the Fermi level leading to very small hole orbits centered at $m$ and $a$ high symmetry points. However, due to their small size and the limited $k$-space grid used for the interpolation, it is difficult to visualize them. Interestingly, these pockets are also not observed in our magnetoresistance measurements, suggesting that their small size is indeed real.
II.2 Magnetoresistance measurements
II.2.1 Measurement configurations
@@
In configuration A, a large magnetoresistance is measured, which saturates to about 1400 $\%$ for certain magnetic field orientations as shown in Fig. 2(a)(ii). The SdH oscillations are appreciable from magnetic fields as low as 5 T. On the other hand, configuration B shows that both non-oscillatory and oscillatory components are suppressed as the magnetic field approaches the c-axis, i.e., the current direction, similar to the results in [21]. In contrast, in configuration C both the oscillatory and non-oscillatory components of magnetoresistance could be observed at all angles, even when the magnetic field was aligned with the c-axis. This configuration thus provides access to additional features of the Fermi surface. Notably, we observe that there is a strong linear contribution to the non-oscillatory magnetoresistance across all field orientations in this case. Moreover, when the magnetic field is close to the c-axis, the non-oscillatory component of the magnetoresistance displayed a secondary linear regime, with a reduced slope. This linear magnetoresistance is accompanied by an oscillatory component with a very large frequency that is absent at other magnetic field orientations. In a subsequent section we will discuss the possible causes of both the linear magnetoresistance in TaTe4 and this high frequency component.
II.2.2 Oscillatory component of the magnetoresistance
@@
The analysis for different frequency regimes was motivated by recent ARPES measurements [20, 16, 18]. These studies revealed that the most intense spectral weight at the Fermi level are ellipsoids centered at the X/M-point in reciprocal space which follow the periodicity of the high-symmetry non-CDW phase, with a cross sectional area about  0.4 $\AA ^{-2}$. According to the Onsager relation, such pocket would correspond to a frequency close to 4000 T when field is perpendicular to the a-c plane, which corresponds to 0° and 90° in configuration A (Figs. 3(a)(i,ii)) and 0° for configurations B and C (Figs. 4(a)(i,ii)). However, no frequency component is observed around the a-direction in the high frequency regime in any configuration. Therefore, there is no evidence of such pocket from the SdH measurements. In configuration A and for this high frequency range, the only contribution occurs in a limited range of field orientations close to 45°, where frequencies of 1387 T $\pm$ 8 T, 2774 T $\pm$ 11 T and 4163 T $\pm$ 15 T are detected (see Fig. 3 (a)(i)). These contributions correspond to the fundamental, first, and second harmonics of the 1387 T signal. Although this frequency value lies in the range of closed orbits that could originate in the FS of the non-CDW phase, the angle dependence of this signal does not agree with that of the predicted non-CDW pockets, which are displayed in Fig. 3(a)(iii).
@@
Since $\beta$ and $\epsilon$ frequencies are the most dominant in the FFT spectrum, it was possible to track their evolution with temperature and calculate their effective masses. To estimate the contribution of charge carriers to the electronic properties of TaTe4, the Liftshitz-Kosevich (LK) formula for the oscillatory magnetoresistance $\tilde{\rho}(B)$, as presented in equation 1 of SI, is used [25]. By fitting the evolution of the oscillatory magnetoresistance with temperature as predicted by this equation, the effective cyclotron masses of each pocket for a fixed angle, $m^{cyc}_{\theta}$, can be obtained. In our case, the effective masses were estimated by filtering the oscillatory signal of each pocket with a band-pass filter to select a single frequency, resulting in a better fit to our data than by performing the fit on the sum of all frequencies with the whole LK formula.
Using this approach and following the magnetic field orientations defined for configuration A in previous sections, we obtained $m^{cyc}_{\beta-0^{\circ}}=(0.30\pm 0.03)m_{e}$ for the lower frequency branch of pocket $\beta$ at 0 and $m^{cyc}_{\beta-45^{\circ}}=(0.32\pm 0.03)m_{e}$ at 45 in configuration A (see SI and Fig. S1 for more details), which suggests that the mass of charge carriers in these pockets is highly isotropic along this rotation plane.
[…]
@@
The $\alpha$ pocket is anisotropic in terms of effective mass and shape. Its charge carriers effective mass smoothly changes with angle, with $m^{cyc}_{\alpha-0^{\circ}}=(0.21\pm 0.03)m_{e}$ at 0 and $m^{cycl}_{\alpha-45^{\circ}}=(0.10\pm 0.02)m_{e}$ at 45 in configuration A. Further details of the effective mass analysis can be found in SI.
@@
In contrast to configuration B, configuration C, for which the magnetic field is always perpendicular to the current, exhibited a clear evolution of different dominant oscillation frequencies at all angles in the aca-c plane, as shown in Figs. 4(a)(i) and (b)(i). In this case, we defined two different frequency regimes for the analysis: a low/mid frequency regime below 2000 T and a high frequency regime, from 2000 T to 5000 T. As in configuration A, the low/mid-frequency regimes dominated the oscillations in configuration C (Fig. 4(a)(i)). The evolution of the $\beta$ pocket, already clearly identified in configuration A, could be tracked at angles near the a-axis; however, its frequency gradually increases and eventually disappears as the magnetic field orientation approaches the c-axis — an effect also observed in configuration B. In contrast, as the $\beta$ pocket oscillation faded, another frequency of oscillation appeared in the range of 1040 T $\pm$ 8 T at 35° and 829 T $\pm$ 7 T at 90°. This frequency follows a clear cylinder-like angular dependence following $1/\cos(\pi/4-\theta)$, in agreement with the DFT predicted $\zeta$ pocket (Figs. 4(a)(ii,iii)). This feature had not been reported in previous magnetotransport studies, likely due to the experimental challenges of imposing current along the a-axis in TaTe4 derived from the quasi-1D morphology of the crystals, as discussed earlier.
II.2.3 High frequency regime and magnetic breakdown effects
@@
The results discussed in the previous two subsections have provided evidence of four out of the six predicted pockets for the CDW-phase FS, with no evidence of contributions coming from bands remanent from the non-CDW high symmetry phase, in contrast to results by recent ARPES works [18, 16, 19]. Therefore we can conclude that the bulk transport band structure of this material is dominated by the CDW-reconstructed bands, which closely match the size and shape of the FS predicted by DFT.
II.2.4 Linear magnetoresistance
@@
III Conclusions
In summary, our combine high-field magnetotransport and DFT study provides a comprehensive reconstruction of the Fermi surface of TaTe4 in its charge density wave (CDW) phase. We resolve four of the theoretically predicted Fermi surface pockets and find no evidence of non-CDW states, confirming that the bulk electronic structure is fully reconstructed by the CDW. Our measurements reveal a previously unobserved quasi-cylindrical pocket and a high-frequency quantum oscillation arising from magnetic breakdown between reconstructed Fermi surface sheets, from which we estimate a CDW gap of approximately 0.29 eV, consistent with ARPES-based estimates.
Beyond the Fermi surface mapping, our results uncover a robust linear magnetoresistance that emerges at all field angles when current flows along the $a$-axis. Notably, a second linear regime appears at high magnetic fields near the $c$-axis, which we associate with magnetic breakdown near hot spots in the Fermi surface. However, further research is needed to fully elucidate whether the linear magnetoresistance, particularly in the low-field regime, can be related to the topological character of this compound. Together, these findings establish TaTe4 as a model platform to investigate the interplay between topological features and correlation-driven band reconstruction. They also open new routes for exploring magnetic breakdown physics, anisotropic transport, and topological responses in quasi-one-dimensional semimetals with strong electronic correlations.
IV Materials and methods
IV.1 Computational details of the electronic structure calculations
[…]
IV.2 Sample growth
IV.3 Sample characterization and transport measurements